Check if (1, 3, 0) lies in the range of a linear
operator T on R3 given by
T (x1, x2 , x3) = (x1 , x2+ x3, x1– x2).
Is T one-one and onto ? Give reasons.
1
Expert's answer
2020-05-21T17:25:15-0400

To check if b=(1,3,0) lies in the range of T:


The corresponding matrix A of the given linear map T is:

"A= \begin{bmatrix} 1 & 0 & 0 \\ 0 & 1 & 1 \\ 1 & -1 & 0 \end{bmatrix}"

the row reduced echelon form of this matrix is:


"M=\begin{bmatrix} 1 & 0 & 0 \\ 0 & 1 & 1 \\ 0 & 0 & 1 \end{bmatrix}" ...(1)


Now, let A(x1,x2,x3)=b = (1,3,0)

Then, x1-x2=0 which implies x1=x2 ,

x2+x3= 3 which implies x3= 2

x1= 1 which implies x2=1

Thus, there exists c=(1,1,2) in R3 such that T(1,1,2)=(1,3,0)

So,(1,3,0) lies in the range of T.


The linear map T is one to one if and only if null space of T has only a trivial solution.

null space of T is given by T(x1,x2,x3) = (0,0,0)

now T(x1,x2,x3) = (0,0,0)

that is A(x1,x2,x3) = (0,0,0)

From (1),

x1=x2=x3 = 0

Since this is a trivial solution, T is one to one.


T is onto:

T is onto if and only if columns of A span R3.


Since the row reduced echelon form of corresponding matrix A of the linear map T in (1) has a pivot in each row,

thus, columns of A are linearly independent, hence, columns of A span R3.


T is onto map.